OpenStudy (anonymous):

A sample of gas has a volume of 0.500 L at a temperature of 27 degrees C and a pressure of 0.4 atm. What is the number of moles in this sample? a) 0.008 b) 0.045 c) 0.091 d) 0.182

OpenStudy (anonymous):

pV = nRT --> n= pV/RT = 0,4 * 101325 Pa * 0,5 * 10-3 m3 / 8,314 J mol-1 K-1 * 27+273 K = 0,008 mol so answer is A)